# OCRBench v2: An Improved Benchmark for Evaluating Large Multimodal Models on Visual Text Localization and Reasoning

### v1

@@ Abstract
Scoring the Optical Character Recognition (OCR) capabilities of Large Multimodal Models (LMMs) has witnessed growing interest recently. Existing benchmarks have highlighted the impressive performance of LMMs in text recognition; however, their abilities in certain challenging tasks, such as text localization, handwritten content extraction, and logical reasoning, remain underexplored. To bridge this gap, we introduce OCRBench v2, a large-scale bilingual text-centric benchmark with currently the most comprehensive set of tasks (4×4$\times$4 × more tasks than the previous multi-scene benchmark OCRBench), the widest coverage of scenarios ($31$ diverse scenarios including street scene, receipt, formula, diagram, and so on), and thorough evaluation metrics, with a total of $10,000$ human-verified question-answering pairs and a high proportion of difficult samples. After carefully benchmarking state-of-the-art LMMs on OCRBench v2, we find that $36$ out of $38$ LMMs score below $50$ ($100$ in total) and suffer from five-type limitations, including less frequently encountered text recognition, fine-grained perception, layout perception, complex element parsing, and logical reasoning. The benchmark and evaluation scripts are available at https://github.com/Yuliang-Liu/MultimodalOCR.
@@ 1 Introduction
Figure 1: Large multimodal models fail to deal with text-intensive tasks accurately. They are prone to errors in tasks such as text localization, handwritten content extraction, and mathematical reasoning, revealing limitations in tackling complex textual information within visual contexts.
@@
The emergence of Large Language Models (LLMs) [1, 101, 8] has greatly improved the understanding and generation of structured text. However, in reality, much of the textual content is unstructured; it appears within images, videos, and other non-textual media in varied positions, orientations, and shapes. The need for processing such unstructured content leads to the study of Large Multimodal Models (LMMs) [5, 53, 139] that extend the text-only LLMs to additional modalities. By pretraining on multimodal data, LMMs acquire the zero-shot ability to interpret across diverse media such as recognizing and understanding complex visual scene text [59]. Such capability represents a significant advancement over standard Optical Character Recognition (OCR), because LMMs not only spot text but also interpret its semantic relevance to a scene.
@@
Compared with classic OCR that typically relies on task-specific models to spot text, the increasing capability of LMMs to process and understand multimodal inputs has opened new potential to redefine the area of OCR. OCR has therefore become an important aspect of recent LMM evaluations. Some text-focused tasks have been included in standard benchmarks to assess the proficiency of LMMs in recognizing and interpreting textual content [26, 121]. Typically, text-based Visual Question Answering (VQA) datasets [93, 7, 107] are repurposed to evaluate OCR by framing generic VQA into questions that require accurate reading of embedded text. However, many of these text-centric datasets are initially created for classic OCR models, which is of limited diversity, depth, and suitability for evaluating LMMs. A common drawback is that, many questions lack sufficient complexity to assess the reasoning abilities of LMMs on scene text, and some can even be answered without visual input [10, 107].
@@
More recently, several preliminary work [58, 46, 103, 50, 42] has attempted to explore the OCR capabilities of LMMs with customized benchmarks. For example, OCRBench [58] consolidates a range of common OCR tasks to evaluate LMM performance across traditional OCR functions. Other datasets, such as ComTQA [132] and ChartX [111], focus on structured text interpretation like table and chart understanding. While such effort represents a leap over standard OCR benchmarks, they remain limited in both data diversity and data quantity (see Table 1), which often leads to rapid performance saturation. For example, recent LMMs such as Qwen2-VL [105] have achieved 96.7% accuracy on the DocVQA dataset [71], nearly matching human performance at 98.1%, and 87.7% on OCRBench [58]. This raises an important question to the OCR community: Have OCR tasks truly been solved in the LMM era, or do existing benchmarks fail to capture the broader challenges present in diverse environments?
@@
To answer the question above, we conducted preliminary tests with several state-of-the-art LMMs, including both open-source models such as Qwen2-VL-8B [105] and Monkey [48], and commercial models like GPT-4o [79]. These tests assessed performance on core OCR tasks, such as text localization, handwritten content extraction, and document-based logical reasoning. As illustrated in Figure 1, each model can fail on one of text-intensive tasks. These failures reveal a gap in detailed visual perception across different models, which constrains their effectiveness in tasks requiring accurate text localization, recognition, and contextual understanding within images. Such observations suggest that existing OCR benchmarks may be insufficient to assess the capabilities and limitations of LMMs for practical OCR applications due to limited task diversity, contextual complexity, and scale.
To bridge this gap, we propose OCRBench v2, a comprehensive benchmark designed to assess LMMs across a diverse set of OCR tasks. As shown in Figure 2, OCRBench v2 assesses eight core text-reading abilities, including text recognition, text referring, text spotting, relation extraction, element parsing, mathematical calculation, visual text understanding, and knowledge reasoning, organized into a total of $23$ concrete tasks. This benchmark provides $10,000$ high-quality, human-validated instruction-response pairs and also six types of evaluation metrics, which offers a rigorous framework for evaluating LMM performance in complex, practical OCR scenarios. In summary, the contributions of this work are three-fold:
  • OCRBench v2: an improved benchmark designed to evaluate the OCR capabilities within the context of LMMs. It assesses eight core OCR competencies and covers $23$ tasks across $31$ diverse scenarios ranging from scientific documents to natural scenes, which provides a thorough evaluation framework encapsulating fundamental and advanced text-centric challenges.
  • We systematically evaluate $38$ representative LMMs, ranging from commercial APIs to open-source models, which establishes broad baselines for OCR performance and enables a comparative understanding of model capabilities across varied OCR tasks.
  • We provide a detailed analysis to identify factors affecting the OCR capabilities of LMMs and highlight key insights for improvement. The analysis examines performance across various dimensions such as model generalization to diverse text types, model robustness, and the ability to tackle complex visual-textual relation.
@@ 2 Related Work
We review work related to OCR-enhanced LMMs and benchmarks for text-centric LMMs.
Figure 3: Sample visualizations for each task in OCRBench v2. The dataset provides a multilingual evaluation setting, comprising 23 sub-tasks grouped under 8 core OCR capabilities. Tasks marked with Refer to caption contain both English and Chinese instructions, while other tasks are either English-only Refer to caption or Chinese-only Refer to caption (Zoomed in for better clarity).
@@
Inspired by the success of LLMs, visual encoders are integrated into LLMs to create LMMs capable of processing both visual and textual information. Early LMMs exhibit strong zero-shot OCR capabilities, which motivates further exploration into text-centric LMMs. One pioneering approach, LLaVAR [131], enhances OCR abilities by training LLaVA [53] on text-centric datasets. mPLUG-DocOwl [118] constructs an instruction-tuning dataset with visual-text understanding to improve the model’s performance. However, these early methods are restricted to low-res inputs, limiting their ability to recognize dense and small text. To address this limitation, attention has shifted to increasing the input resolution of LMMs. DocPedia [25] processes high-res images in the frequency domain without increasing the input sequence length. UReader [119] crops an image into multiple sub-images, thereby achieving fine-grained perception of high-res text. On the contrary, LayoutLLM [67] adopts the LayoutLMv3 [33] as the vision encoder to process the high-res images. As the resolution of inputs increases, so do the challenges related to computational efficiency. To tackle this issue, TextMonkey [59] proposes a Token Resampler to compress redundant visual feature tokens, while mPLUG-DocOwl $1.5$ [31] introduces an H-Reducer to reduce the number of image tokens. As these models advance, they have achieved remarkable results on established benchmarks. Nevertheless, challenges remain unsolved in certain key areas such as text localization, entity extraction, and logical reasoning.
@@
To evaluate LMMs, developing a comprehensive benchmark is essential. Previous effort has focused on creating scenario-specific benchmarks to assess LMMs in particular contexts. For example, DocVQA [71] is designed to evaluate the document comprehension abilities of LMMs, while ChartQA [69] is tailored to chart interpretation skills. Similarly, Infographics VQA [73] is dedicated to assessing the understanding of infographic images. Additionally, TextVQA [93] aims to evaluate text comprehension in real-world scenes. To further investigate the robustness of the model, some methods expand the scope of evaluation scenarios. OCRBench [58] introduces a holistic evaluation framework that covers five core OCR tasks. CONTEXTUAL [103] is developed with context-sensitive instructions. SEED-Bench-2-Plus [46] encompasses a wide spectrum of text-rich images from various sources, including web content, maps, and charts. To provide a more thorough assessment, some benchmarks design multiple evaluation tasks within a specific scenario. TableVQA-Bench [42] first focuses on VQA tasks in the table domain. MMTab [134] and ComTQA [132] then extend the task scope, including table detection, structure recognition, and table querying. Moreover, ChartY [9], ChartX [111], and MMC [51] evaluate LMMs in chart understanding through tasks such as chart information extraction and reasoning. In this work, we focus on establishing a new benchmark called OCRBench v2, which contains more tasks than previous benchmarks and provides a systematic evaluation framework to reveal the limitations of LMMs in diverse text-rich environments.
3 OCRBench v2 Benchmark
In this section, we introduce OCRBench v2 from four aspects, including task, annotation, statistics, and evaluation.
@@ 3.1 Task Description
To provide a comprehensive evaluation framework for text-reading tasks, we categorize OCR capabilities into eight core areas, each encompassing specific sub-tasks that address various aspects of text comprehension and interpretation. Figure 3 demonstrates representative examples for each task, which showcases both visual inputs and corresponding instructions. Detailed descriptions of these core capabilities are as follows.
@@
Text Referring. Determining the location of texts is necessary for real-world OCR applications. OCRBench v2 evaluates this ability with text grounding and VQA with position tasks, requiring LMMs to localize text positions accurately.
@@
Element Parsing. Apart from text recognition, LMMs face the need of parsing complex elements for downstream applications. This ability is evaluated via table parsing, chart parsing, document parsing, and formula recognition.
Mathematical Calculation. Mathematical calculation abilities are essential for LMMs to address tasks that require numerical reasoning. Hence, text counting is introduced to assess the textual perception ability by counting the number of text instances. Additionally, we enhance the original math QA data by rendering textual questions into images, accompanied by geometric figures.
Visual Text Understanding. To tackle sophisticated tasks involving human interaction, LMMs need to comprehend semantic information of texts, a capability we term visual text understanding. This ability is evaluated by document classification and diagram QA. Additionally, we include basic VQA instructions where answers are located directly within the image, which refers to cognition VQA.
Knowledge Reasoning. Certain tasks require reasoning based on world knowledge and involve complex inference procedures, including science QA, APP agent interactions, ASCII art classification, and text translation. Further, we collect some complicated VQA instructions, where answers are not directly visible in images, termed reasoning VQA.
Due to space limitations, additional details for each subtask can be found in Appendix C.
Figure 4: The ratios of each type of tasks in OCRBench v2.
@@ 3.2 Annotation Curation
In this section, we introduce annotation curation in three parts, including dataset collection, instruction formulation, and manual verification.
Dataset Collection. To curate diverse data for OCRBench v2, we manually harvest and screen $81$ text-rich academic datasets. More details about the data sources can be found in Appendix B. To ensure diverse scenarios coverage, we also supplement them with additional private data. In all, our dataset comprises $31$ typical scenarios [^1] The $31$ scenarios include schematic diagrams, scientific papers, text image patches, filled tables, charts, receipts, question contexts, mathematical formulas, product labels, phone screenshots, indoor scenes, industry research reports, posters, street scenes, ASCII Art, shop signs, financial reports, chemical formulas, textbooks, magazines, emails, web screenshots, details page, verification code, resumes, illustrations, newspapers, road signs, menus, notify, and questionnaires., and the number of images in each scene is displayed in Table 9.
Instruction Formatting. To convert existing annotations into the LMM-compatible instruction format, we design specific prompts for each task. For complex tasks such as document parsing that require structured output, we include a format example to minimize the impact of instruction-following ability and focus the evaluation on OCR capabilities. Additionally, considering the distinct training strategies for localization tasks in LMMs, we standardize the coordinates by normalizing them with image sizes and by scaling to the range of $[0,1000]$. Such a standardization is explicitly specified in the related prompts. For unlabeled data, we manually annotate them according to task requirements.
Manual Verification. To ensure data quality, we manually review all instructions and correct approximately $1\%$ annotation errors.
Figure 5: OCR-related statistics of OCRBench v2.
Figure 6: Comparison of average entropy, type-token ratio, and variability index between OCRBench v2 and OCRBench.

### v2

@@ Abstract
Scoring the Optical Character Recognition (OCR) capabilities of Large Multimodal Models (LMMs) has witnessed growing interest. Existing benchmarks have highlighted the impressive performance of LMMs in text recognition; however, their abilities in certain challenging tasks, such as text localization, handwritten content extraction, and logical reasoning, remain underexplored. To bridge this gap, we introduce OCRBench v2, a large-scale bilingual text-centric benchmark with currently the most comprehensive set of tasks (4×4$\times$4 × more tasks than the previous multi-scene benchmark OCRBench), the widest coverage of scenarios ($31$ diverse scenarios), and thorough evaluation metrics, with $10,000$ human-verified question-answering pairs and a high proportion of difficult samples. Moreover, we construct a private test set with $1,500$ manually annotated images. The consistent evaluation trends observed across both public and private test sets validate the OCRBench v2’s reliability. After carefully benchmarking state-of-the-art LMMs, we find that most LMMs score below $50$ ($100$ in total) and suffer from five-type limitations, including less frequently encountered text recognition, fine-grained perception, layout perception, complex element parsing, and logical reasoning. The project website is at: https://99franklin.github.io/ocrbench_v2.
@@ 1 Introduction
The emergence of Large Language Models (LLMs) [1, 2, 3] has greatly improved the understanding and generation of structured text. However, in reality, much of the textual content is unstructured; it appears within images, videos, and other non-textual media in varied positions, orientations, and shapes. The need for processing such unstructured content leads to the study of Large Multimodal Models (LMMs) [4, 5, 6] that extend the text-only LLMs to additional modalities. By pretraining on multimodal data, LMMs acquire the zero-shot ability to interpret across diverse media, such as recognizing and understanding complex visual scene text [7]. Such capability represents a significant advancement over standard Optical Character Recognition (OCR), because LMMs not only spot text but also interpret its semantic relevance to a scene.
Figure 1: Large multimodal models struggle with text-intensive tasks accurately. They are prone to errors in tasks like text localization, handwritten content extraction, and mathematical reasoning, revealing limitations in tackling complex textual information within images.
@@
Compared with classic OCR that typically relies on task-specific models to spot text, the increasing capability of LMMs to process multimodal inputs has opened new potential to redefine the area of OCR. OCR has therefore become an important aspect of recent LMM evaluations. Some text-focused tasks have been included in standard benchmarks to assess the proficiency of LMMs in recognizing and interpreting textual content [8, 9]. Typically, text-based Visual Question Answering (VQA) datasets [10, 11, 12] are repurposed to evaluate OCR by framing generic VQA into questions that require accurate reading of embedded text. However, many of these datasets are initially created for classic OCR models, which are of limited diversity, depth, and suitability for evaluating LMMs. A common drawback is that, many questions lack sufficient complexity to assess the reasoning abilities of LMMs on scene text, and some can even be answered without visual input [13, 12].
@@
More recently, several customized benchmarks [14, 15, 16, 17, 18] have explored the OCR capabilities of LMMs. For example, OCRBench [14] consolidates $5$ core text-oriented tasks to evaluate LMM performance across traditional OCR functions. Other datasets, such as ComTQA [19] and ChartX [20], focus on structured text interpretation like table and chart understanding. While such effort represents a leap over standard OCR benchmarks, they remain limited in both data diversity and quantity (see Tab. 1), often leading to rapid performance saturation. For example, recent LMMs such as Qwen2.5-VL [21] have achieved 96.4% accuracy on the DocVQA dataset [22], nearly matching human performance at 98.1%, and 88.8% on OCRBench [14]. This raises an important question for the community: Do models perform well enough on text-oriented visual understanding tasks in the LMM era, or do existing benchmarks fail to capture the broader challenges in diverse environments?
To answer the question above, we conducted preliminary tests with several state-of-the-art LMMs, including Qwen2.5-VL-7B [21], InternVL3-14B [23], and GPT-4o [24]. These tests assessed performance on text-oriented tasks, such as text localization, handwritten content extraction, and document-based logical reasoning. As illustrated in Fig. 1, each model can fail on one of the text-intensive tasks. These failures reveal a gap in detailed visual perception across different models, which constrains their effectiveness in tasks requiring accurate text localization, recognition, and contextual understanding within images. Recent benchmarks, such as OmniDocBench [25], CC-OCR [26], and MMLONGBENCH-DOC [27], have broadened evaluation to cover more comprehensive scenarios, including fine-grained document parsing and multi-page document understanding. Their analyses reveal the limited capabilities of LMMs for practical OCR applications and highlight the growing need for benchmarks that allow for more robust and varied evaluation of LMMs.
To bridge this gap, we propose OCRBench v2, a comprehensive benchmark designed to assess LMMs across diverse text-oriented visual understanding tasks. As shown in Fig. 3, OCRBench v2 assesses eight core text-reading abilities, including text recognition, text referring, text spotting, relation extraction, element parsing, mathematical calculation, visual text understanding, and knowledge reasoning, organized into a total of $23$ concrete tasks. This benchmark provides $10,000$ high-quality, human-validated instruction-response pairs and also six types of evaluation metrics, which offers a rigorous framework for evaluating LMM performance in complex, practical OCR scenarios. For better evaluation quality, we further collect and label $1,500$ additional text-images from scratch, reserved as the private test set. This private data serves as an independently curated test set to validate model generalization. In summary, the contributions of this work are three-fold:
  • OCRBench v2: an improved benchmark designed to assess eight core OCR competencies and covers $23$ tasks across $31$ diverse scenarios, which provides a thorough evaluation framework encapsulating fundamental and advanced text-centric challenges.
  • We systematically evaluate state-of-the-art LMMs, ranging from commercial APIs to open-source models, which establishes broad baselines for OCR performance and enables a comparative understanding of model capabilities across varied text-oriented visual understanding tasks.
  • We provide a detailed analysis to identify factors affecting the OCR capabilities of LMMs. The analysis examines performance across various dimensions such as model generalization to diverse text types, model robustness, and the ability to tackle complex visual-textual relations.
@@ 2 Related Work
Inspired by LLMs, visual encoders are integrated into them to create LMMs capable of processing both images and text. Early LMMs exhibit strong zero-shot OCR capabilities, motivating the exploration of text-centric LMMs. For instance, some work [30, 31] use text-centric instruction-tuning to enhance OCR-related abilities. But they are restricted to low-res inputs, limiting the ability to recognize dense and small text. To address this, several studies [32, 33, 34] shift attention to increasing the input resolution. As the resolution of inputs increases, so does computational cost. To tackle this issue, TextMonkey [7] introduces a Token Resampler to compress redundant visual feature tokens, mPLUG-DocOwl2 [35] presents a DocCompressor module for compressing high-res images, and DocKylin [36] adopts adaptive pixel slimming and dynamic token slimming modules to reduce redundant regions. To enhance layout perception, DocLayLLM [37] integrates layout information into LMMs inputs, LayTokenLLM [38] shares position IDs between text and layout tokens, DocMark [39] utilizes adaptive generation of markup languages to build structured document representations, while Marten [40] introduces an additional mask generator during pre-training. Despite strong results on existing benchmarks, challenges remain unsolved in certain key areas such as text localization, entity extraction, and logical reasoning.
@@
Previous efforts have focused on creating scenario-specific benchmarks to assess LMMs. For example, DocVQA [22], ChartQA [41], Infographics VQA [42], and TextVQA [10] evaluate models on document understanding, chart reasoning, infographic interpretation, and scene text comprehension, respectively. To broaden evaluation scope, OCRBench [14] introduces a holistic evaluation framework covering five text-oriented tasks, while CONTEXTUAL [16] and SEED-Bench-2-Plus [15] introduce context-sensitive and diverse real-world images. Other benchmarks target specific challenges such as dense text understanding [43], complex structure parsing [26], and fine-grained document analysis [25]. To provide a more thorough assessment, some benchmarks design multiple tasks within a specific scenario. TableVQA-Bench [18], MMTab [28], and ComTQA [19] explore table-based tasks, while ChartY [44], ChartX [20], and MMC [29] focus on chart information extraction and reasoning. Recently, DUDE [45], MM-NIAH [46], MP-DocVQA [47], MMLONGBENCH-DOC [27], and LongDocURL [48] explore the long document understanding capability of LMMs. In this work, we establish OCRBench v2, a systematic benchmark to reveal the limitations of LMMs in diverse single-image, text-related scenarios.
3 Why Do We Need OCRBench v2?
Limitations of Existing Benchmarks.
Recent evaluations of LMMs’ OCR capabilities have made significant progress, yet most existing benchmarks exhibit limitations. Datasets like DocVQA, ChartQA, and TextVQA are often narrow in scope, focusing predominantly on text recognition within specific domains such as forms, tables, or documents. While useful for isolated capabilities, they fall short in task diversity, instruction complexity, and structured output formats that better reflect the multimodal nature of LMMs. In particular, many of these benchmarks were originally tailored for traditional OCR systems that prior to the emergence of LMMs. Furthermore, as illustrated in Fig. 2, complex task-specific processes are needed for LMMs when extended to more text-oriented tasks, which limits the evaluation of their broader capabilities.
Figure 2: As evaluation for LMMs expands to diverse text-oriented tasks, existing datasets often require task-specific handling, making unified and scalable evaluation difficult.
The Necessity of Unified Multi-task Evaluation.
With the emergence of LMMs, current models show generalization ability to handle multiple tasks. To assess these multi-task models, unified benchmarks like LongDocURL [48], OmniDocBench [25], CCOCR [26], OCRBench [14], CONTEXTUAL [16], SEED-Bench-2-Plus [15], have been proposed and successfully demonstrated the value of evaluating text-oriented models across diverse tasks. These benchmarks show the importance of unified evaluation frameworks in guiding model development. However, as model capabilities expand, existing benchmarks with limited task coverage result in fragmented and sometimes misleading insights. To address this, a unified benchmark is essential to: 1) Understand generalization: Can a model perform consistently across varied text-centric tasks? 2) Diagnose failure models: Does a model that excels in recognition also succeed in reasoning, localization, and parsing? 3) Guide model development: Unified evaluation provides clearer signals for architecture and training improvements.
As shown in Fig. 3, OCRBench v2 tackles this by combining 23 tasks under 8 core capabilities within one framework. This holistic design enables systematic comparison of models and highlights trade-offs (e.g., performance on reasoning vs. recognition) that isolated benchmarks cannot reveal.
How OCRBench v2 Addresses the Gaps.
OCRBench v2 is a comprehensive, and high-difficulty benchmark specifically built to evaluate LMMs in realistic OCR settings, with key advantages: 1) Breadth of coverage: With 31 scenarios, we ensure diverse contextual challenges; 2) Task variety: The benchmark spans 8 OCR-related capabilities, many of which are poorly handled by current LMMs; 3) Instruction complexity: Human-authored prompts and structured outputs (e.g., Markdown, JSON, LaTeX) raise the bar beyond simple answer extraction; 4) Private evaluation test set: To prevent overfitting and training contamination, we additionally provide a private test set.
Ultimately, OCRBench v2 fills a critical gap by offering a unified and challenging benchmark that reflects the practical needs of OCR in the LMM era. It not only measures what current models can do, but more importantly, reveals what they still cannot.
Design Rationale: Focusing on Single-Image Text Tasks.
While designing OCRBench v2, we focus on challenges in single-image, text-related scenarios, and do not extend our study to multi-image tasks. This design choice is grounded in two considerations: 1) Single-image understanding is the foundation for more complex multimodal tasks. Many existing models still perform unsatisfactorily in various single-image scenarios, which motivates our work; 2) Given long-context inputs, multi-page tasks have more emphasis on long-sequence modeling, requiring specific benchmarks to assess this capability individually. For example, MMLONGBENCH-DOC focuses on evaluating the ability of LMMs to locate and understand content across pages in long documents.
Private Dataset for Reliable Evaluation.
To further enhance the assessment quality, we also construct a private test set. This data comprises $1,500$ manually collected text-rich images with human-annotated labels, covering 23 tasks aligned with the distribution of the public data. The images are from diverse sources, including printed books, e-books, scanned documents, and web content. During data collection and annotation, we meticulously curated samples to align with practical text-oriented applications. Given that benchmarks may be contaminated in massive internet-scraped pre-training data of LMMs, this data will not be released. Instead, we maintain a regularly updated leaderboard to reflect the performance of advanced LMMs. Moreover, consistent performance trends and model rankings observed on both the public and private test sets (see Section 5.2) indicate the benchmark’s well-founded design and its effectiveness in identifying model capabilities.
4 Benchmark Construction
Figure 3: Sample visualizations for each task. OCRBench v2 comprises 23 sub-tasks grouped under 8 core OCR capabilities. Tasks marked with Refer to caption contain both English and Chinese instructions, while other tasks are either English-only Refer to caption or Chinese-only Refer to caption (Zoomed in for better clarity).
In this section, we describe the task description, annotation curation, statistics, and evaluation criteria. Due to space limitations, more details can be found in the Appendix.
@@ 4.1 Task Description
To provide a comprehensive evaluation framework for text-reading tasks, we categorize OCR capabilities into eight core areas, each encompassing specific sub-tasks that address various aspects of text comprehension and interpretation. Fig. 3 exhibits samples for each task, with visual inputs and corresponding instructions. Detailed descriptions of these core capabilities are as follows.
@@
Text Referring. Determining the location of texts accurately is necessary for real-world OCR applications. This ability is evaluated with text grounding and VQA with position tasks.
@@
Element Parsing. LMMs face the need of parsing complex elements for downstream applications. This ability is evaluated via table parsing, chart parsing, document parsing, and formula recognition.
Mathematical Calculation. Math calculation is essential for LMMs to address numerical reasoning tasks. Hence, text counting is introduced to assess the textual perception ability. Besides, we enhance the math QA data by rendering textual questions into images, accompanied by geometric figures.
Visual Text Understanding. To tackle sophisticated tasks involving human interaction, LMMs need to comprehend the semantic information of texts, a capability we term visual text understanding. This ability is evaluated by document classification and diagram QA. Additionally, we include basic VQA instructions where answers are located directly within the image, which refers to cognition VQA.
Knowledge Reasoning. Some tasks require complex inference and world knowledge, including science QA, APP agent interactions, ASCII art classification, text translation, and reasoning VQA (where answers are not directly visible in images).
@@ 4.2 Annotation Curation
Dataset Collection. To ensure data diversity, we manually harvest and screen $81$ text-rich academic datasets. To ensure diverse scenario coverage, we also supplement them with additional private data. In all, our dataset comprises $31$ typical scenarios (see Appendix for the full list).
Instruction Formatting. To convert existing annotations into the LMM-compatible instruction format, we design specific prompts for each task. For complex tasks such as document parsing that require structured output, we include a format example to minimize the impact of instruction-following ability and focus the evaluation on OCR capabilities. Additionally, considering the distinct training strategies for localization tasks in LMMs, we standardize the coordinates by normalizing them with image sizes and by scaling to the range of $[0,1000]$. Such a standardization is explicitly specified in the related prompts. For the completeness of the tasks and scenarios, we carefully annotate the additional images to ensure they align with the designed task requirements.
Manual Verification. To ensure data quality, we manually review all instructions for public data and correct approximately $1\%$ annotation errors.
Figure 4: OCR-related statistics and prompt quality assessment of OCRBench v2.
4.3 Statistics of OCRBench v2
Here we present the OCR-related statistics and the measurement of prompt quality. As shown in Fig. 4 (a) and (b), we count the distribution of line-level OCR results of $7,400$ English and $2,600$ Chinese images. And Fig. 4 (c) exhibits the average number of line-level OCR results per category. These statistics demonstrate that the text information is sufficiently rich in OCRBench v2. In addition, Fig. 4 (d) compares the Average Entropy, Type-Token Ratio, and Average Variability Index of the questions between OCRBench v2 and OCRBench. OCRBench v2 presents higher values across all three metrics, indicating more diverse, less redundant, and structurally varied questions. This suggests it provides a more comprehensive and challenging benchmark for LMMs.